OpenStudy (anonymous):

You borrowed $5,000 from your parents to purchase a used car. You have agreed to make payments of $250 per month plus an additional 1% interest on the unpaid balance of the loan. a. Is this problem an example of a geometric series or an arithmetic series? Support your answer mathematically by applying the concepts from this unit. b. Find the first year’s monthly payments that you will make and the unpaid balance after each month. c. Find the total amount of interest paid over the term of the loan.

OpenStudy (amistre64):

Does that mean that the 1% is compounded monthly?

OpenStudy (anonymous):

That's kinda what it sounds like to me....

OpenStudy (amistre64):

first payment of 250 + 5000(.01) = 300; remaining balance of 4750?

OpenStudy (anonymous):

So 300 a month? I can figure the different left over balances after each month I think. How would I find the total amoun of interest paid just add up the percentages until I reach zero for my payments?

OpenStudy (amistre64):

thats the long way to do it :) set up payment coupons of 250 + unpaid(.1)...but I think this is just a loan that is compounded monthly, just not too sure :)

OpenStudy (amistre64):

5000/250 would be the number of times needed to pay the principal off

OpenStudy (anonymous):

oh ok so it would take 20 payments?

OpenStudy (amistre64):

at 20 payments the principal is gone :)

OpenStudy (amistre64):

5000(.01) + (5000-250)(.01) + (5000 - 500)(.01).... for 20 times

OpenStudy (amistre64):

(5000 - n(250))(.01)? does that sound right?

OpenStudy (anonymous):

Ok so this is sorta making sense.... when you say for 20 times do I multiply all of that by 20?

OpenStudy (anonymous):

n being 20 right?

OpenStudy (amistre64):

lol.... I wish I knew fer sure :) We know it takes 20 payments to pay that loan off; and each payment made subtracts 250 from the principal..

OpenStudy (anonymous):

right

OpenStudy (amistre64):

so there is a relation here that happens 20 times....

OpenStudy (anonymous):

ok I understand that

OpenStudy (amistre64):

im glad one of us does :)

OpenStudy (anonymous):

lol well I think I do anyways sorta ha ha

OpenStudy (amistre64):

first iteration is: (5000 - 0(250))(.01) = interest paid second iteration: (5000 - 1(250))(.01) = interest paid third iteration: (5000 - 2(250))(.01) = interset paid .....

OpenStudy (anonymous):

oh ok I was trying to but 20 in for n and was getting 0 which I knew wasn't right

OpenStudy (amistre64):

lets clean this up: x = 5000 ; m=250 ; t = .01 xt - mt(n-1) = interest paid in any given iteration...how do we add these all up?

OpenStudy (anonymous):

so the first interest paid is 50 second interest paid is 47.50 third interest paid is 45

OpenStudy (anonymous):

hold on here we are trying to find the answer to question c right

OpenStudy (amistre64):

good, good..... theres an answer to find? :)

OpenStudy (amistre64):

yes..c = total interest paid

OpenStudy (amistre64):

if we can determine how much interest is paid for any given number of months, we will know the answer to "b"

OpenStudy (anonymous):

ohhh ok so I have to find the interest paid for 12 months

OpenStudy (amistre64):

yes.... and there is an equation you can use, just trying to figure it out :)

OpenStudy (amistre64):

(xt - mt(0)) + (xt - mt(1)) + (xt - mt(2)) + (xt - mt(3))... xt - mt(0) + xt - mt(1) + xt - mt(2) + xt - mt(3).... what can we factor out of these?

OpenStudy (amistre64):

xt (-mt(0)- mt(1) - mt(2) - mt(3) -mt(4))...

OpenStudy (anonymous):

oh gosh I'm terrible at factoring

OpenStudy (amistre64):

xt (-mt (n-1)) looks to be it; for any given number of months... x = 5000 ; m=250 ; t = .01 50 (-2.5 (n-1)) sounds right

OpenStudy (amistre64):

not quite there.... at n=1 we get 0 for interest....

OpenStudy (amistre64):

whats the total interest we know for the first 3?

OpenStudy (amistre64):

142.50 is the first 3 months of interest: 50 (-2.5 (3-1)) = 250 .... so it aint right yet :)

OpenStudy (amistre64):

xt - mt(0) + xt - mt(1) + xt - mt(2) + xt - mt(3).... xt(1 - mt(n-1) + 1 mt(n-1) + 1 - mt(n-1))...

OpenStudy (amistre64):

100 (1 - 250(n-1)) when n=3 nope.... i guess we need someone smarter than me :)

OpenStudy (amistre64):

2.5(21-n) will give you the amount of interest paid on a certain iteration..... so far, now just need to integrate that with respect to n ...maybe :)

OpenStudy (amistre64):

<html> <body> number of iterations: <input id="inp1"><br><br> <input type = button value ="Calculate Total interest" onClick="qwe()"><br><br> <input id="inp2" value="total"><br><br> Amount of interest this payment:<input id="inp3"> </body> <script language=javascript> var j, i=0, t=0, x,c function qwe() { x = document.getElementById("inp1").value x=x*1 for (j=1;j<x+1;j=j+1) { i=2.5*(21-j) t=t+i } document.getElementById("inp2").value=t document.getElementById("inp3").value=i t=0 } </script> </html> Thatll do it :)

OpenStudy (amistre64):

c) is $525 of interest is paid at the end of the loan (20 months). b): jan: 250+ 50 = 300 ; 4750 feb: 250+47.50 = 297.50 ; 4500 mar: 250+45 = 295 ; 4250 apr: 250+42.50 = 292.50 ; 4000 may: 250+40 = 290 ; 3750 jun: 250+37.50 = 287.50 ; 3500 jul: 250+35 = 285 ; 3250 aug: 250+32.50 = 282.50 ;3000 sep :250+30 = 280 ; 2750 nov: 250+27.50 = 277.50 ;2500 dec: 250+25 =275 ; 2250

OpenStudy (amistre64):

a) id guess arithmetical series....

OpenStudy (anonymous):

missing october in part b.... If it's arithmetic though I think I just do Oct:250+27.50 = 277.50 ;2500 Nov:250+25 =275 ; 2250 Dec:250+22.50=272.5; 2000